you are given two vectors:

v=[x2 +y2+ z2, 2xyz, x+y+2z]

u=[xy+z , xy2 z2 , x+3z]

Calculate the following expressions:

a) v+u

b) v•u

c) v x u

d) div v
Answer #1

Here ts-( 2,2x47, 2+4+22) I22,22+32) (a) 0+1=(++24ytZ,2792+で8 td1221132 tu cy+z, x242+ 2+6 +5z) AD +292y3234 +32) t 22) =yt22)(x +2122) (d) dhv 220+2xz+2 div V= 2+2+2 x2 A